Blog

Day five of IOP 2.  Weather fine and calm, scattered cloud later.

Four soundings on usual IOP schedule.  Trained Steve Oncley with the 9:15am sounding (see photo), he learnt well!

Air conditioner output jumps around, seems to vary between about 60F and 70F even though we're keeping the setting steady on 67F.   The profiler power fluctuates between about 4.7 and 5.2 kW as a result.  As of 19 UT the levels were 5.18kW, 34.03 dBm, 50V, 11.7 A.

Did a stock-take of supplies:as of midday there were 66 sondes in trailer, 56 balloons, 14 small and 2 large Helium tanks remaining.  Apparently there is more Helium at Kemp and more sondes in storage.  Checked on Helium usage - 10 tanks used to launch 97 sondes so getting an average of 9.7 sondes per tank.  However that's a mix of small and large tanks, so it appears we're getting about 11 soundings from the large tanks and about 8 from the small tanks.

Also filled in more of the ISS instrument spreadsheet (serial numbers, locations etc).

Day four of IOP2.   Weather mainly cloudy with a uniform stratus deck most of the day and a light northeast wind.

Four soundings following the regular IOP schedule during the day, plus an extra sounding last night around 4UT for comparisons with the profiler and lidar.

The trailer AC temperature fluctuated during day resulting in fluctuating power from the 449 power amplifier. When the ambient temperature rose to around 22C, the power out dropped to about 4.75kW, but when the ambient temperature dropped to 18C the amp power was 5.07kW.  Discovered the AC was on a schedule, so removed that and put on a fixed setting.  Additional fans have been ordered for the amp.

Cleared out filter on sodar AC and tipping bucket rain gauge (note that brief 15.24 mm/hr reading at 20:30 UTC was due to this maintenance and should be ignored).  Kyle's last day and one of his last jobs was cleaning some bird poop from the Allsky camera around 15 UT - thanks Kyle!

Brian being photographed by a reporter doing the 22 UT sounding with Ankur observing.

Day three of IOP2.  Weather fine, more morning mist than previous, then clear, then scattered clouds developing from mid-morning along with a light northerly breeze.

Four regular IOP soundings with a plan for an extra sounding tonight to check on how the profilers are doing at night.  Visit by reporters from the Milwaukee Journal Sentinel who recorded the 1pm sounding.

Profiler: working well. The power was a little low this morning at 4.7kW, but this was likely because the AC had been turned up overnight, leading to a temperature of 30C just inside the amp enclosure on the rack.  Turning the AC back down and adding a fan blowing into the front of the enclosure lowered the enclosure temperature to 25C (cf ambient of 20C)  and boosted to power back to 5kW.   Readings as of midday: power 5.02kW 34.03 dBm, 50V 11.52A.  The King Air was briefly seen in the spectral display during a close flyover at 10:10am (1510UT) and likely is showing up briefly during other fly-bys.

Cleaned off windranger lidar window (had trace of soot) and Allsky camera dome (lots of dew) at 13 UT.

Dan doing a survey of the ISS site using the Leica scanner laser from a USGS survey marker located on the north edge of our field. 

Day two of IOP 2.  Weather clear early, then clouds developed and a couple of lines of rain passed mid-morning.  Clouds remained for the afternoon.

Four soundings as usual for the IOP, although Brian requested the 9:15am sounding be launched a few minutes since the King Air accelerated it's schedule to avoid the mid-morning rain.

Eckhard reported that the sodar-RASS cross-talk has been a little high, so we rotated the receiver antenna about 20 degrees (around 1830UT) and now the cross-talk is at a better level (around 2V instead of 4V).  The RASS range does seem to be reaching a little higher this afternoon as compared to recent afternoons.  The snake was hanging out on the Trimble on the shelf above the sodar desk, then along the top of the door.  Kyle corralled it outside but it'll probably come back.

The photo shows a King Air fly-by at 10am (15UT), just prior to the first of the rain showers. 

IOP2 began today with a sounding at 6:10am (1110 UTC), followed by soundings at 9:15am (1315 UT), 1pm (18 UT) and 4:45pm (2145 UT). This will be the sounding schedule for the rest of the week.

All soundings and other systems working well.  Kyle cleaned NOAA sensors and I cleaned ceilometer, both cameras and WS800 solar sensor around 13 UT.  Also cleaned trailer AC filter which was very clogged with dust.

The King Air did a short test flight and NOAA operated their drones at the WLEF tower.  

The weather was clear (apart from a little early mist) and calm with a light south-westerly from morning.  

Profiler still running in non-over sampling mode and reporting winds to around the 300m level.  Power 5 kW, 34.03 dBm, 50V, 11.66A.  

We did a series of delay line tests to attempt to understand the relatively high lowest wind measurements we've been seeing.  The lowest reported winds have been around 400m or higher.  The delay line with patch antenna was positioned on channel 1 from about 1835 UT until about 1945 UT - also later in day from about 2255 - 2335 UT.  Tested various modes including varying pulse codes (1,2,4) and first gate settings.  Eventually found that we get the lowest winds with the oversampling off - not sure why as oversampling appeared to work okay at Savant.  Anyway with it off we now get winds down to 325m (had been 400m or higher).  Examination of profiles of parameters such as fading time and auto-correlation offset with the delay line in place hint that the RX delay may set set a little high (see attached image - ignore the winds at 0 height).  Will look at this more closely in post-processing, but it appears we may be able to adjust the lowest winds down to around 200 - 230 m.


All systems operating well.  Weather raining overnight and this morning, with westerly breeze. fortunately clearing around midday.  Regular sounding at 18UT.

We hosted a CHEESEHEAD open house this afternoon, with an extra sounding at 2134 UT and about 30 members of the public in attendance.


Profiler: A little water accumulated on the antenna (just a millimeter or 2). Cleared channel 1 off about 16:20 UT, then the other channels around 16:30 UT - seemed to make little difference to the signal.  The profiler computer was very sluggish so rebooted that in the process.  See a separate blog post for some adjustments lowest range gate and delay line testing.  The snake was hanging out under the 449 antenna today.




Update 17 August 2019

All systems running well. Weather fine and warm, scattered cloud, light variable winds.  

Profiler power 5kW, 34 dBm, 50V 11.5A.  Experimented with adjusting profiler lowest range gate since we don't seem to be getting wind measurements down at the lowest gate, or often even at 400m level.  Lowest level was 300m, raised to 400 around 2030UT, then dropped to 350m around 2115UT.  Now getting more consistent winds at 400m, however this still seems too high.  Will experiment with delay line tomorrow. Did 18Z sounding and some data backups, both to disk and DVDs.  Visited some isfs sites and KIT trailer at tower.  Didn't find the sodar trailer snake but did see a black cat.

Update 08/16/2019

Soundings have been going off fine. Profiler reading nominal values, all instruments have been up since the last update. Internet went out a couple days ago for approximately 21 hours but then came back online and no data was lost in the process. The backup hard drive has been replaced for a fresh one (CHEESEHEAD_3). The snake is reading nominal values. 

Update 08/13/2019

All systems functioning as expected. Wisconsin taking over sounding duties today for three days. Sounding went off a little late (1:08) due to the sounding computer software having to be restarted, but on restart everything was functional. Balloon popped at about 100hPa which is a lower than nominal. One of the hard drives is filling up and about at maximum capacity, Bill said he would let me know when to change it out. 

Quite a bit of rain came through last night, but did not effect any instruments. 

Around 2:00 PM Central time (~19:00 UTC) the network connections for ISS went down. An email was sent to Bill to see what we should do about it. Around 9:00 PM Central time (~2:00 UTC), after receiving some information from Bill,  Liz and I both came back to the trailer to implement Bill's suggestions of cycling power on the the network switches and cradlepoint. The first several attempts at cycling the power of the networks switches and cradlepoint did nothing. We then attempted to cycle only the cradlepoint one last time (using the reset button on it) and at the same time went around and pushed the physical ethernet connections into their seats to ensure they were physically connected. The network came back online and the majority of errors in Nagios cleared up. This morning, all stray Nagios errors have cleared up and ISS looks good (Note: there are still a few warnings, but they don't look to be impeding any operations). Liz says she contacted Bill again and was told that the instrument data from the time of the network outage is available, but will not be processed automatically. 

It also looks like the ISFS trailer may have had a short power outage last night as the microwave was flashing 12:00 and a battery charging power supply was reset. However we are not sure if this outage also hit the ISS trailer. The wind profiler and all other ISS instruments are running properly as if no power outage had occurred. 


Snake update: We are hoping the snake is still in the Sodar-Rass trailer, it got quite cold and wet outside last night and may have not been safe for the snake to be outside through the night. We will manually check on the status of the snake later today.

ISS Ops 8/9/2019

All Iss instruments are up and data is looking good. Kyle arrived last night. Profiler power readings are: 5 kW, 34.03 dBm, and 50V at 11.5A. Checked up on Snake, and it is still happily residing in the Sodar-Rass trailer.

SODAR-RASS Repair


After receiving instruction on how to shut down/ power up the system I changed the SSM2142 amplifier chip in the RASS receiver box (picture shown below). The system is back up and running and just waiting for data to come in to verify if this fixed the issue. I do also believe we have a resident snake in the SODAR trailer as you can see in the second picture. 




Today there was a high of 79F and skies were partly cloudy. All ISS instruments are running well except for the RASS. Working with Bill and Eckhard through email communication to troubleshoot and hopefully get up and running again soon. Eckhard suggested changing out an amplifier (part number SSM2142) as long as we had spare parts. I checked the trailer and was able to find many extra of these. Now just waiting further instruction on changing these out. Today I assisted Steve in changing batteries at the SW sites as well as troubleshooting a TRH probe at SW3 and installed soil sensors as SW3 too. Sounding went smoothly with a burst at 97.31 hPa. The profiler power levels read at: 5.01 kW, 34.03 dBm, and 50V @ 11.51A.

Liz arrived Saturday night and had first day at project sight today. All ISS instruments are up and looking good. Changed batteries at NW1, NW2, and NW4 with Steve in the morning. Brian helped with sounding in the afternoon and went over NOAA sensor dome cleaning as well.